7kg mass is being pulled toward the right with a rope that has a tension of 63 newtons at an angle of 42 degrees to the x-axis if the surface is frictionless what acceleration will the mess develop?
1
Expert's answer
2018-10-17T15:43:09-0400

The force in the x-direction is

F=T cos⁡α.

The acceleration will be

a=F/m=T/m cos⁡α.

a=63/7 cos⁡42=6.7 m/s^2 .